OpenStudy (anonymous):

Find the midpoint of the line segment whose endpoints are (-2, 5) and (4, -9).

OpenStudy (anonymous):

to find a midpoint you do the following take the two x-points (here they're -2 and 4) and add them. then divide the sum by two so that would be (-2+4)/2 = 2/2 =1 this is your x coordinate take the two y coordinates (here they're 5 and -9) and add them, then divide by tow. so here (5-9)/2 = (-4)/2 = -2 this is the y cooridnate so you get (1, -2)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

anytime :)